# Spin-dependent terms of the Breit-Pauli Hamiltonian evaluated with explicitly correlated Gaussian basis set for molecular computations

### v1

@@ Abstract
This work compiles the spin-dependent leading-order relativistic and quantum electrodynamical corrections to the electronic structure of atoms and molecules within the non-relativistic quantum electrodynamics framework. We report the computation of perturbative corrections using an explicitly correlated Gaussian basis set, which allows high-precision computations for few-electron systems. In addition to numerical tests for triplet Be, triplet $\mathrm{H}_{2}$, and triplet $\mathrm{H}_{3}^{+}$ states and comparison with no-pair Dirac–Coulomb–Breit Hamiltonian energies, numerical results are reported for electronically excited states of the triplet helium dimer, He2, for which the present implementation delivers high-precision magnetic coupling curves necessary for a quantitative understanding of the fine structure of its high-resolution rovibronic spectrum.
@@ I Introduction
For few-particle systems, an explicitly correlated basis set is frequently applied in combination with a variational procedure to capture the essential electron-electron correlation effects, leading to fast convergence in energy and other physical properties [9, 10]. To obtain analytic expressions for the matrix elements, floating Explicitly Correlated Gaussians (fECGs) are used as a spatial basis set [9, 10, 11, 12, 13], for which numerical efficiency in the high-precision computation of few-electron systems has already been demonstrated [14, 15, 16, 17, 18, 19, 20, 21].
@@
Since the electronic energy of few-particle systems can be converged to high-precision in a variational fECG approach, the relativistic and quantum electrodynamical (QED) contributions become ‘visible’ and important even for (compounds of) light elements. High-resolution spectroscopic measurements reveal small magnetic effects, seen as fine and hyperfine splittings in the spectra. We will refer to the relevant interactions as ‘spin-dependent’ relativistic (and QED) interactions, and we will also use the term ‘spin-independent’ relativistic and QED corrections, which give contributions to the centroid energy. While the computation of the centroid corrections is well elaborated, including the regularization techniques [22, 23, 24], which enhance the convergence of singular terms in the Gaussian basis representation (for the missing incorrect electron-nucleus and electron-electron coalescence behaviour).
@@
In this paper, we focus on the fECG-based implementation of the spin-dependent matrix elements of the Breit–Pauli Hamiltonian. We also include corrections for the anomalous magnetic moment of the electron, which constitute the leading-order ($\alpha^{3}E_{\text{h}}$, with the $\alpha$ fine structure constant) QED corrections to the spin-dependent terms. We note that higher-order QED ($\alpha^{4}E_{\text{h}}$) corrections have been derived by Douglas and Kroll in 1974 for the helium atom triplet states [25] and more recently, the $\alpha^{5}E_{\text{h}}$ order corrections non-radiative, radiative and recoil corrections have been derived and evaluated for triplet helium states [18, 26]. Inner shell transitions are in excellent agreement with experiment, though there is a significant deviation for the ionization energy between theory [27, 28] and experiment [29, 30, 31, 32].
@@
The present work focuses on the relativistic leading-order QED corrections, i.e., up to $\alpha^{3}E_{\text{h}}$ order, but goes beyond atomic applications; molecular systems require the development of a floating ECG methodolgy, reported in the present work.
Non-floating, spherically symmetric, ECG basis sets have already been successfully used in spin-dependent atomic computations, considering beryllium, boron and carbon atomic states [33, 34, 35, 36, 19, 37].
@@
In this paper, we report a general $N$-particle implementation without restrictions on the number and spatial arrangement of the atomic nuclei. The implementation is based on a spinor basis representation, in which the spatial and spin components are evaluated independently and subsequently combined via calculating simple tensor products. The approach is, in principle, ‘general’ and not restricted by the number of particles or point-group symmetry. For testing the developed methodology, we perform computations for triplet $\mathrm{Be}$ for comparison with literature values. Two-electron triplet $\mathrm{H}_{2}$ and also two-electron triplet $\mathrm{H}_{3}^{+}$ (both a linear and a triangular configuration) are compared with our in-house developed no-pair DCB computations [38].
Then, as a first application of the methodology, we compute magnetic coupling curves for several electronically excited states of the triplet He2 molecule, which significantly improve upon previously available values using standard quantum chemistry methodologies [39, 40, 41, 42, 43, 44, 45]. These newly-computed coupling curves are used in rovibronic-fine-structure computations reported in separate papers [46, 47].
@@ II Theoretical framework
For atoms and molecules with low nuclear charge, the non-relativistic description often provides a good starting point. The non-relativistic electronic energy is incremented with relativistic and QED corrections arranged according to powers of the fine-structure constant, $\alpha$. In this paper, we include all terms up to $\alpha^{3}$ (in Hartree atomic units),
@@
$\hat{H}^{(0)}|\varphi^{(0)}\rangle=E^{(0)}|\varphi^{(0)}\rangle\ ,$ (2)
where the non-relativistic electronic Hamiltonian is
$\hat{H}^{(0)}$ $=-\frac{\hbar^{2}}{2m}\sum_{i=1}^{n_{\text{el}}}\boldsymbol{\Delta}_{\boldsymbol{r}_{i}}-\sum_{i=1}^{n_{\text{el}}}\sum_{A=1}^{N_{\text{nuc}}}\frac{Z_{A}}{r_{iA}}+\sum_{i=1}^{n_{\text{el}}}\sum_{j>i}^{n_{\text{el}}}\frac{1}{r_{ij}}\;,$ (3)
with the nuclei clamped at $\boldsymbol{R}_{A}$. We use the short notation, $r_{iX}=|\boldsymbol{r}_{iX}|$, $\boldsymbol{r}_{iA}=\boldsymbol{r}_{i}-\boldsymbol{R}_{A}$, and $\boldsymbol{r}_{ij}=\boldsymbol{r}_{i}-\boldsymbol{r}_{j}$.
The leading-order relativistic corrections, $\alpha^{2}E^{(2)}$, can be computed using the non-relativistic wave function, $\varphi^{(0)}$, as the expectation value of the Breit-Pauli (BP) Hamiltonian. The focus of this work is the computation of the spin-dependent contributions, so we write the Hamiltonian as the sum of spin-independent (sn) and spin-dependent (sd) terms [48, 49]
$\hat{H}^{(2)}=\hat{H}^{(2)}_{\text{sn}}+\hat{H}^{(2)}_{\text{sd}}=\hat{H}_{\text{MV}}+\hat{H}_{\text{D1}}+\hat{H}_{\text{OO}}+\hat{H}_{\text{D2}}+\hat{H}_{\text{SO}}+\hat{H}_{\text{SOO}}+\hat{H}_{\text{SOO'}}++\hat{H}_{\text{SOO}}+\hat{H}_{\text{SS}}\ ,$ (4)
where the manifestly spin-independent terms are re-iterated only for completeness, i.e., the mass-velocity (MV), the Darwin one (D1), the Darwin two (D2), and the orbit-orbit terms (OO) are
@@
$\hat{H}_{\text{OO}}=-\frac{1}{2}\sum_{i=1}^{n_{\text{el}}}\sum_{j>i}^{n_{\text{el}}}\left[\frac{1}{r_{ij}}\hat{\boldsymbol{p}}_{i}\hat{\boldsymbol{p}}_{j}+\frac{1}{r_{ij}^{3}}(\boldsymbol{r}_{ij}(\boldsymbol{r}_{ij}\hat{\boldsymbol{p}}_{j})\hat{\boldsymbol{p}}_{i})\right]\ .$ (6)
The corrections carrying spin operators in their expressions are the spin-orbit interaction, with $\hat{\boldsymbol{l}}_{iX}=\boldsymbol{r}_{iX}\times\hat{\boldsymbol{p}}_{i}$ and $\hat{\boldsymbol{s}}_{i}=\frac{1}{2}I(1)\otimes...\otimes\boldsymbol{\sigma}(i)\otimes...\otimes I(n_{\text{el}})$ with the $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ Pauli matrices,
@@
which is written as the sum of the magnetic dipole-dipole interaction,
@@
and the Fermi contact term,
@@
In addition, we consider the effect of the anomalous magnetic moment of the electron [50, 51, 52] on the spin-dependent terms, giving rise to the leading-order QED correction for the spin-dependent part,
@@ III Implementation
$|\varphi^{(\Gamma,SM_{S})}\rangle=\sum_{\mu=1}^{N}C_{\mu}|\phi^{(\Gamma,SM_{S})}_{\mu}\rangle\ ,$ (17)
where the coefficients $C_{\mu}$ are obtained by diagonalizing the non-relativistic Hamiltonian matrix. We define the basis function as
$\phi^{(\Gamma,SM_{S})}_{\mu}=\phi^{(\Gamma,SM_{S})}(\boldsymbol{r};\boldsymbol{A}_{\mu},\boldsymbol{s}_{\mu},\boldsymbol{\theta}_{\mu})=\hat{\mathcal{A}}\{f^{(\Gamma)}(\boldsymbol{r};\boldsymbol{A}_{\mu},\boldsymbol{s}_{\mu})\chi^{(SM_{S})}(\boldsymbol{\theta}_{\mu})\}\ .$ (18)
$\chi^{(SM_{S})}(\boldsymbol{\theta}_{\mu})$ is a spin function, $f^{(\Gamma)}(\boldsymbol{r};\boldsymbol{A}_{\mu},\boldsymbol{s}_{\mu})$ is a spatial function. In this work, it is the symmetrized floating explicitly correlated Gaussian function (fECG) [9, 10, 12, 11, 13]
@@
where $\hat{\mathcal{P}}_{\Gamma}$ is the projector onto the $\Gamma$ irreducible representation (irrep) of the point-group defined by the fixed nuclear skeleton, $\boldsymbol{I}_{3}$ is the three-dimensional identity matrix, $\boldsymbol{A}_{\mu}$ and $\boldsymbol{s}_{\mu}$ are parameters to be optimized by minimization of the non-relativistic energy.
@@
Regarding the spin function $\chi^{(SM_{S})}_{\mu}=\chi^{(SM_{S})}(\boldsymbol{\theta}_{\mu})$, $S$ is the total electron spin quantum number and $M_{S}$ is the quantum number for the spin projection. $S,M_{S}$ and the $nel$ define a degenerate spin space in which the spin functions can be expressed by a linear combination (with a linear so-called $\theta$-parameterization) of the elementary spin functions (Supporting Information). The norm of the parameter vector is fixed to 1, thereby reducing the number of free parameters of the spin space. In the non-relativistic energy minimization, we also include the $\boldsymbol{\theta}_{\mu}$ parameter vector.
@@
$=\sum_{\mu=1}^{N}\sum_{\nu=1}^{N^{\prime}}C^{\ast}_{\mu}C^{\prime}_{\nu}\langle\hat{\mathcal{A}}\{f_{\mu}\chi_{\mu}\}|\hat{\mathcal{H}}|\hat{\mathcal{A}}\{f^{\prime}_{\nu}\chi^{\prime}_{\nu}\}\rangle$
@@
$\text{SOO}:\quad\hat{G}^{\text{SOO}}_{ia}$ $=-\frac{1}{2}\sum_{j>i}^{n_{\mathrm{el}}}\frac{1}{r_{ij}^{3}}\hat{{l}}_{ij,a}\ ,$ (24)
$\text{SOO}^{\prime}:\quad\hat{G}^{\text{SOO}^{\prime}}_{ia}$ $=-\sum_{j>i}^{n_{\mathrm{el}}}\frac{1}{r_{ij}^{3}}\hat{{l}}_{ji,a}\ ,$ (25)
@@
$\langle\varphi|\hat{H}_{\text{SS,dp}}|\varphi^{\prime}\rangle$ $=\sum_{\mu=1}^{N}\sum_{\nu=1}^{N^{\prime}}C^{\ast}_{\mu}C^{\prime}_{\nu}\sum_{r=1}^{N_{\text{perm}}}(-1)^{p_{r}}\sum_{i,j=1}^{n_{\text{el}}}\sum_{a,b=1}^{3}\langle f_{\mu}|\hat{G}_{ij,ab}|\hat{P}_{r}f^{\prime}_{\nu}\rangle_{\boldsymbol{r}}\langle\chi_{\mu}|\hat{s}_{ia}\hat{s}_{jb}|\hat{Q}_{r}\chi^{\prime}_{\nu}\rangle_{\boldsymbol{\sigma}}\;.$ (27)
The spatial integrals are calculated analytically by exploiting the mathematical properties of fECGs, similarly to our previous work, e.g., Ref. [17, 53, 54, 55, 20].
Since the present work is about the evaluation of the spin-dependent Breit-Pauli matrix elements, some notes regarding the spinor structure and matrix elements are appropriate. To evaluate the spin matrix elements, the spin-adapted ‘$\boldsymbol{\theta}$’ parametrization (see also Refs. 56 and 46) is transformed into the spinor (vector) representation (Supporting Information),
$|\chi^{(SM_{S})}(\boldsymbol{\theta})\rangle=\sum_{\begin{subarray}{c}m_{s_{1}},m_{s_{2}},...,m_{s_{n_{\text{el}}}}\\ \Sigma m_{s_{i}}=M_{S}\end{subarray}}d_{m_{s_{1}}m_{s_{2}}...m_{s_{n_{\text{el}}}}}^{(SM_{S})}|\Xi_{m_{s_{1}}m_{s_{2}}...m_{s_{n_{\text{el}}}}}\rangle\ ,$ (28)
where $|\Xi_{m_{s_{1}}m_{s_{2}}...m_{s_{n_{\text{el}}}}}\rangle=|\eta_{m_{s_{1}}}\rangle\otimes|\eta_{m_{s_{2}}}\rangle\otimes\dots\otimes|\eta_{m_{s_{n_{\mathrm{el}}}}}\rangle$, and $|\eta_{s_{i}}\rangle$ is a single-particle spin function (vector). Depending on the $m_{s_{i}}=-\frac{1}{2}$ or $+\frac{1}{2}$ value, the spin function is ‘spin up’, $|\eta_{0}\rangle=|\alpha\rangle=\left(\begin{array}[]{c}1\\ 0\end{array}\right)$, or ‘spin down’, $|\eta_{1}\rangle=|\beta\rangle=\left(\begin{array}[]{c}0\\ 1\end{array}\right)$.
Then, the $\hat{\boldsymbol{s}}_{i}$ spin operator is considered as the tensor product of identity matrices and the Pauli matrix,
@@
Then, the matrix elements of the spin functions is written as (where the tensor structure can be further exploited)
@@
In a nutshell, the following algorithm has been implemented in our in-house developed fECG-based computer program, named QUANTEN, to compute matrix elements of spin-dependent relativistic operators connecting various electronic-spin states of small molecules (and also to atoms, of course):
@@
    The spin function $\chi_{\mu}^{(SM_{S})}$ with parameters $\boldsymbol{\theta}_{\mu}$ are transformed to the spinor (vector) representation, Eq. (28), to obtain the $d_{\boldsymbol{a},\mu}^{(SM_{S})}$ coefficients.
@@
    The spin operators (matrices) $\boldsymbol{s}_{i}$ are constructed according to Eq. (29), and the ladder operators (matrices) are generated,
    $\boldsymbol{s}^{\pm}_{i}$ $=\boldsymbol{s}_{i,x}\pm\text{i}\boldsymbol{s}_{i,y}\ ,\quad\boldsymbol{S}^{\pm}=\sum_{i=1}^{n_{\mathrm{el}}}\boldsymbol{s}^{\pm}_{i}\ ,$ (34)
@@
    The spatial and spin contributions are combined according to Eqs. (22) and (27), to construct the relativistic (and QED) corrections (coupling matrix elements of the electron-spin states).

### v2

# Spin-dependent terms of the Breit-Pauli Hamiltonian evaluated with an explicitly correlated Gaussian basis set for molecular computations

@@ Abstract
This work collects the spin-dependent leading-order relativistic and quantum-electrodynamical corrections for the electronic structure of atoms and molecules within the non-relativistic quantum electrodynamics framework. We report the computation of perturbative corrections using an explicitly correlated Gaussian basis set, which allows high-precision computations for few-electron systems. In addition to numerical tests for triplet Be, triplet $\mathrm{H}_{2}$, and triplet $\mathrm{H}_{3}^{+}$ states and comparison with no-pair Dirac–Coulomb–Breit Hamiltonian energies, numerical results are reported for electronically excited states of the helium dimer, He2, for which the present implementation delivers high-precision magnetic coupling curves necessary for a quantitative understanding of the fine structure of its high-resolution rovibronic spectrum.
@@ I Introduction
For few-particle systems, an explicitly correlated basis set is frequently used in combination with a variational procedure to capture the essential electron-electron correlation effects, leading to fast convergence in the energy and other physical properties [9, 10]. To obtain analytic expressions for the matrix elements, floating Explicitly Correlated Gaussians (fECGs) are used as a spatial basis set [9, 10, 11, 12, 13], for which numerical efficiency in the high-precision computation of few-electron systems has already been demonstrated [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Since the electronic energy of few-particle systems can be converged to high precision in a variational fECG approach, the relativistic and quantum electrodynamical (QED) contributions become ‘visible’ and important even for (compounds of) light elements. High-resolution spectroscopic measurements reveal small magnetic effects, seen as fine and hyperfine splittings in the spectra. We will refer to the relevant interactions as ‘spin-dependent’ relativistic (and QED) interactions, and we will also use the term ‘spin-independent’ relativistic and QED corrections, which contribute to the centroid energy (defined as the average of the fine-structure energy levels corresponding to the degenerate non-relativistic subspace). While the computation of the centroid corrections is well elaborated, including the regularisation techniques [28, 29, 30], which enhance the convergence of singular terms in the Gaussian basis representation (for the incorrect electron-nucleus and electron-electron coalescence behaviour).
@@
In this paper, we focus on an fECG-based implementation of the spin-dependent matrix elements of the Breit-Pauli Hamiltonian. We also include corrections for the anomalous magnetic moment of the electron, which constitutes the leading-order ($\alpha^{3}E_{\text{h}}$, with the $\alpha$ fine-structure constant) QED corrections to the spin-dependent terms. We note that higher-order QED ($\alpha^{4}E_{\text{h}}$) corrections have been derived by Douglas and Kroll in 1974 for the helium atom triplet states [31] and more recently, the $\alpha^{5}E_{\text{h}}$-order corrections have been derived and evaluated for triplet helium states [18, 32]. Inner shell transitions are in excellent agreement with experiment, though there is a significant deviation for the ionization energy between theory [33, 34] and experiment [35, 36, 37, 38].
@@
The present work focuses on the relativistic and leading-order QED corrections, i.e., up to $\alpha^{3}E_{\text{h}}$ order, but goes beyond atomic applications; molecular systems with clamped nuclei require the development of a floating ECG methodology, reported in the present work.
[…]
@@
In this paper, we report a general $N$-electron implementation without restrictions on the number and spatial arrangement of the clamped atomic nuclei. The implementation is based on a spinor basis representation, in which the spatial and spin components are evaluated independently and subsequently combined via calculating simple tensor products. The approach is, in principle, ‘general’ and not restricted by the number of electrons or point-group symmetry. For testing the developed methodology and computer implementation, we perform computations for triplet $\mathrm{Be}$ for comparison with literature values. As to molecular systems (the main target of this work), two-electron triplet $\mathrm{H}_{2}$ and also two-electron triplet $\mathrm{H}_{3}^{+}$ (both a linear and a triangular configuration) are also included in the test set. For molecular systems, high-accuracy benchmark values are not available in the literature; so, we tested these results with our in-house developed no-pair Dirac–Coulomb–Breit methodology (currently available for two-spin-1/2-fermion systems) [48], in which the spin-dependent Breit-Pauli relativistic effects appear at lowest order of the fine-structure constant. Then, as a first large-scale application of the implementation, motivated by recent and ongoing experiments,[49, 50, 51] we compute magnetic coupling curves for several electronically excited states of the (triplet and singlet) He2 molecule, which significantly improve upon previously available quantum chemistry computations [52, 53, 54, 55, 56, 57, 58]. The newly-computed coupling curves are used in rovibronic-fine-structure computations reported in separate papers [26, 27].
@@ II Theoretical framework
For atoms and molecules with low nuclear charge, the non-relativistic description often provides a good starting point. The non-relativistic electronic energy is incremented with relativistic and QED corrections arranged according to powers of the $\alpha$ fine-structure constant. In this paper, we include all terms up to $\alpha^{3}$ (in Hartree atomic units),
@@
$\hat{H}^{(0)}|\varphi^{(0)}\rangle=E^{(0)}|\varphi^{(0)}\rangle$ (2)
for the non-relativistic electronic Hamiltonian,
$\hat{H}^{(0)}$ $=-\frac{1}{2}\sum_{i=1}^{n_{\text{el}}}\boldsymbol{\Delta}_{\boldsymbol{r}_{i}}-\sum_{i=1}^{n_{\text{el}}}\sum_{A=1}^{N_{\text{nuc}}}\frac{Z_{A}}{r_{iA}}+\sum_{i=1}^{n_{\text{el}}}\sum_{j>i}^{n_{\text{el}}}\frac{1}{r_{ij}}\;,$ (3)
with the nuclei clamped at the $\boldsymbol{R}_{A}\ (A=1,\ldots,N_{\text{nuc}})$ positions. We use the short notation, $r_{iX}=|\boldsymbol{r}_{iX}|$, where $X$ is either a nucleus or an electron index, i.e., resulting in $\boldsymbol{r}_{iA}=\boldsymbol{r}_{i}-\boldsymbol{R}_{A}$ or $\boldsymbol{r}_{ij}=\boldsymbol{r}_{i}-\boldsymbol{r}_{j}$.
The leading-order relativistic corrections, $\alpha^{2}E^{(2)}$, can be computed as the expectation value of the Breit-Pauli (BP) Hamiltonian with the non-relativistic wave function, $\varphi^{(0)}$. The focus of this work is the computation of the spin-dependent contributions, so we write the Hamiltonian as the sum of spin-independent (sn) and spin-dependent (sd) terms [59, 60]
$\hat{H}^{(2)}=\hat{H}^{(2)}_{\text{sn}}+\hat{H}^{(2)}_{\text{sd}}=\hat{H}_{\text{MV}}+\hat{H}_{\text{D1}}+\hat{H}_{\text{OO}}+\hat{H}_{\text{D2}}+\hat{H}_{\text{SO}}+\hat{H}_{\text{SOO}}+\hat{H}_{\text{SOO'}}+\hat{H}_{\text{SS}}\ ,$ (4)
where the manifestly spin-independent terms are re-iterated only for completeness, i.e., the mass-velocity (MV), the one-electorn Darwin (D1), the two-electron Darwin (D2), and the orbit-orbit terms (OO) are
@@
$\hat{H}_{\text{OO}}=-\frac{1}{2}\sum_{i=1}^{n_{\text{el}}}\sum_{j>i}^{n_{\text{el}}}\left[\frac{1}{r_{ij}}\hat{\boldsymbol{p}}_{i}\hat{\boldsymbol{p}}_{j}+\frac{1}{r_{ij}^{3}}(\boldsymbol{r}_{ij}(\boldsymbol{r}_{ij}\hat{\boldsymbol{p}}_{j})\hat{\boldsymbol{p}}_{i})\right]\ ,$ (6)
respectively. The corrections carrying spin operators in their expressions are the spin-orbit interaction, with $\hat{\boldsymbol{l}}_{iX}=\boldsymbol{r}_{iX}\times\hat{\boldsymbol{p}}_{i}$ and $\hat{\boldsymbol{s}}_{i}=\frac{1}{2}I(1)\otimes...\otimes\boldsymbol{\sigma}(i)\otimes...\otimes I(n_{\text{el}})$ with the $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ Pauli matrices,
@@
which is written as the sum of the magnetic spin dipole-dipole interaction,
@@
and the Fermi contact term for each pair of electrons,
@@
In addition, we consider the effect of the anomalous magnetic moment of the electron [61, 62, 63] on the spin-dependent terms, which gives rise to the leading-order QED correction for the spin-dependent part (a brief summary with references is in the SM of Ref. 26),
@@ III Implementation
$|\varphi^{(\Gamma,SM_{S})}\rangle=\sum_{\mu=1}^{N}C_{\mu}|\phi^{(\Gamma,SM_{S})}_{\mu}\rangle\$, (17)
where the coefficients $C_{\mu}$ are obtained by diagonalising the non-relativistic Hamiltonian matrix. We define the basis function as
[…]
$\chi^{(SM_{S})}(\boldsymbol{\theta}_{\mu})$ is a spin function, $f^{(\Gamma)}(\boldsymbol{r};\boldsymbol{A}_{\mu},\boldsymbol{s}_{\mu})$ is a spatial function. In this work,symmetrized floating explicitly correlated Gaussian functions (fECG) [9, 10, 12, 11, 13] are used as spatial functions,
@@
where $\hat{\mathcal{P}}_{\Gamma}$ is the projector onto the $\Gamma$ irreducible representation (irrep) of the point-group defined by the fixed nuclear skeleton (technical details are reported in some detail in the Supplementary Material of Ref. 48), $\boldsymbol{I}_{3}$ is the three-dimensional identity matrix, $\boldsymbol{A}_{\mu}$ and $\boldsymbol{s}_{\mu}$ are non-linear parameters, which are generated and optimised by minimisation of the non-relativistic energy [9, 64, 65, 10, 26].
@@
Regarding the spin function $\chi^{(SM_{S})}_{\mu}=\chi^{(SM_{S})}(\boldsymbol{\theta}_{\mu})$, $S$ is the total electron spin quantum number and $M_{S}$ is the quantum number for the spin projection. $S,M_{S}$ for $n_{\text{el}}$ electrons define a spin subspace, for which the spin functions can be expressed as a linear combination of the elementary spin functions; we call this representation the $\theta$-parameterization[9, 64] (see also Supporting Information). In the non-relativistic energy minimization, we also include the $\boldsymbol{\theta}_{\mu}$ parameter vector.
@@
[…]
@@
$\text{SOO}:\quad\hat{G}^{\text{SOO}}_{ia}$ $=-\frac{1}{4}\sum_{j\neq i}^{n_{\mathrm{el}}}\frac{1}{r_{ij}^{3}}\hat{{l}}_{ij,a}\ ,$ (24)
$\text{SOO}^{\prime}:\quad\hat{G}^{\text{SOO}^{\prime}}_{ia}$ $=-\frac{1}{2}\sum_{j\neq i}^{n_{\mathrm{el}}}\frac{1}{r_{ij}^{3}}\hat{{l}}_{ji,a}\ ,$ (25)
@@
$\langle\varphi|\hat{H}_{\text{SS,dp}}|\varphi^{\prime}\rangle$ $=\sum_{\mu=1}^{N}\sum_{\nu=1}^{N^{\prime}}C^{\ast}_{\mu}C^{\prime}_{\nu}\sum_{r=1}^{N_{\text{perm}}}(-1)^{p_{r}}\sum_{i=1}^{n_{\text{el}}}\sum_{j>i}^{n_{\text{el}}}\sum_{a,b=1}^{3}\langle f_{\mu}|\hat{G}_{ij,ab}|\hat{P}_{r}f^{\prime}_{\nu}\rangle_{\boldsymbol{r}}\langle\chi_{\mu}|\hat{s}_{ia}\hat{s}_{jb}|\hat{Q}_{r}\chi^{\prime}_{\nu}\rangle_{\boldsymbol{\sigma}}\;,$ (27)
where
$\hat{G}_{ij,ab}=\frac{\delta_{ab}}{r_{ij}^{3}}-\frac{3r_{ij,a}r_{ij,b}}{r_{ij}^{5}}\ .$ (28)
The spatial integrals are calculated analytically by exploiting the mathematical properties of fECGs, similarly to our previous work, e.g., Ref. 17, 66, 67, 68, 20.
Since the present work is about the evaluation of the spin-dependent Breit-Pauli matrix elements, some notes regarding the spinor structure and matrix elements are appropriate. To evaluate the spin matrix elements, the spin-adapted $\theta$-parametrization (see also Refs. 25 and 26) is transformed into the spinor (vector) representation (see also Supporting Information),
$|\chi_{\mu}^{(SM_{S})}(\boldsymbol{\theta})\rangle=\sum_{\begin{subarray}{c}t_{1},t_{2},...,t_{{n_{\text{el}}}}\\ \Sigma t_{i}=2M_{S}+1\end{subarray}}d_{t_{1}t_{2}...t_{n_{\text{el}}},\mu}^{(SM_{S})}|\Xi_{t_{1}t_{2}...t_{{n_{\text{el}}}}}\rangle\ ,$ (29)
where $t_{i}=2m_{s_{i}}+1$, $|\Xi_{t_{1}t_{2}...t_{n_{\text{el}}}}\rangle=|\eta_{t_{1}}\rangle\otimes|\eta_{t_{2}}\rangle\otimes\dots\otimes|\eta_{t_{n_{\text{el}}}}\rangle$, and $|\eta_{t_{i}}\rangle$ is a single-particle spin function (vector). Depending on the $m_{s_{i}}=+\frac{1}{2}$ or $\frac{1}{2}$ value, the spin function is ‘spin up’, $|\eta_{1}\rangle=|\alpha\rangle=\left(\begin{array}[]{c}1\\ 0\end{array}\right)$, or ‘spin down’, $|\eta_{0}\rangle=|\beta\rangle=\left(\begin{array}[]{c}0\\ 1\end{array}\right)$.
Then, the $\hat{\boldsymbol{s}}_{i}$ spin operator is represented as the tensor product of (two-by-two) identity matrices and the Pauli matrix,
@@
Then, the matrix elements of the spin functions are written as (where the tensor structure can be further exploited)
@@
In a nutshell, the following algorithm has been implemented in our in-house developed fECG-based computer program, named QUANTEN, to compute matrix elements of spin-dependent relativistic operators connecting various electronic-spin states of small molecules (and also atoms, of course):
@@
    The spin function $\chi_{\mu}^{(SM_{S})}$ with parameters $\boldsymbol{\theta}_{\mu}$ are transformed to the spinor (vector) representation, Eq. (29), to obtain the $d_{t_{1}t_{2}...t_{n_{\text{el}}},\mu}^{(SM_{S})}$ coefficients.
@@
    The spin operators (matrices) $\boldsymbol{s}_{i}$ are constructed according to Eq. (30), and the ladder operators (matrices) are generated as
    $\boldsymbol{s}^{\pm}_{i}$ $=\boldsymbol{s}_{i,x}\pm\text{i}\boldsymbol{s}_{i,y}\ ,\quad\boldsymbol{S}^{\pm}=\sum_{i=1}^{n_{\mathrm{el}}}\boldsymbol{s}^{\pm}_{i}\ .$ (35)
@@
    The spatial and spin contributions are combined according to Eqs. (22) and (27), to obtain the relativistic (and QED) coupling matrix elements of the electron-spin states.